OpenStudy (anonymous):

Find all the zeroes of the equation. –3x^4+ 27x^2 + 1200 = 0

OpenStudy (mayankdevnani):

-12+54+1200=0

OpenStudy (mayankdevnani):

1254-12=0

OpenStudy (anonymous):

that's all i have to do?

OpenStudy (anonymous):

is it supposed to be -3x^4 +27x^2+1200=0?

OpenStudy (anonymous):

fixed it. thanks

OpenStudy (mayankdevnani):

oh!!!

OpenStudy (mayankdevnani):

-3x^4+26x^2=-1200

OpenStudy (anonymous):

so then what? is that all i do? this stuff is so confusing :(

OpenStudy (mayankdevnani):

x^2(-3x^2+27)=-1200

OpenStudy (anonymous):

but how do i find "the zeroes" ??

OpenStudy (anonymous):

\[-3x ^{4}+27x ^{2}+1200=0\] \[t=x ^{2}\] \[-3t ^{2}+27t+1200=0\] find the zeroes of t, then substitute them in\[t=x ^{2}\] to get the zeroes you're looking for. Don't forget its always \[\pm \sqrt{t}\]

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

no problem!